OpenStudy (anonymous):

Justin has 6 bags of dog food. His dogs eat 3/4 bag each day. How long will his dog food last?

OpenStudy (danjs):

dividing 6 by 3/4 is the same as multiplying 6 by the flipped over fraction 6 / (3/4) = 6 *(4/3) = 24 / 3 = 8

OpenStudy (danjs):

The good will last 8 days

OpenStudy (danjs):

food